# Geometry-Aware Global Feature Aggregation for Real-Time Indirect Illumination

### v1

@@ 5.3 Validation
Effectiveness of global geometry-aware feature aggregation
@@
As discussed in Sec. 3, we exploit the modified multi-head attention mechanism for guiding global feature aggregation based on geometry features. This allows modeling long-range dependencies for any two locations in the image space, which is inherently impractical for fully convolutional networks (FCN) that prior works used. We validate this with a simple scene containing one major indirect light source (see Fig. 7), whose horizontal span exceeds the receptive field of the compared models. We show that previous works have difficulty capturing distant indirect light sources due to the locality of convolution operations, whereas our method efficiently models long-range dependencies with a feature aggregation module. Please refer to our supplemental video for more comparisons under dynamic lighting.
@@
Moreover, to justify the potential performance loss of the monochromatic design (introduced by the parallel inference of three channels, see Sec. 3), we implement a mini version of the monochromatic network Ours (mini), using the same network architecture as Ours’s but reducing the number of channels in intermediate layers to roughly half the model size. As shown in the first section of Tab. 2, the Ours (mini) still performs better than Normal with similar computational cost and a much smaller amount of parameters. As discussed in Sec. 1, we believe the success of the monochromatic network may result from three major reasons: 1) eliminating the redundant connections (correlations) between different color channels; 2) reducing the high-dimensional space of RGB input/output by introducing the prior that the rendering process of each color channel (or wavelength) is independent; thus 3) making the network more robust to the inequality and bias in the color distribution of training data. To further verify the superiority of our monochromatic model, we introduce scenes illuminated with colored lights that are not present in our training data (see Fig. Geometry-Aware Global Feature Aggregation for Real-Time Indirect Illumination and Fig. 12). Our monochromatic model can still produce plausible outputs without further training or fine-tuning, whereas the normal models generally cannot achieve this with no (or only a few) training data in this domain.

### v2

@@ 5.3 Validation
Effectiveness of global geometry-aware feature aggregation As discussed in Sec. 3, we exploit the modified multi-head attention mechanism for guiding global feature aggregation based on geometry features. This allows modeling long-range dependencies for any two locations in the image space, which is inherently impractical for fully convolutional networks (FCN) that prior works used. We validate this with a simple scene containing one major indirect light source (see Fig. 8), whose horizontal span exceeds the receptive field of the compared models. We show that previous works have difficulty capturing distant indirect light sources due to the locality of convolution operations, whereas our method efficiently models long-range dependencies with a feature aggregation module. Please refer to our supplemental video for more comparisons under dynamic lighting.
@@
Moreover, to justify the potential performance loss of the monochromatic design (introduced by the parallel inference of three channels, see Sec. 3), we implement a mini version of the monochromatic network Ours (mini), using the same network architecture as Ours’s but reducing the number of channels in intermediate layers to roughly half the model size. As shown in the first section of Tab. 2, the Ours (mini) still performs better than Normal with similar computational cost and a much smaller amount of parameters. As discussed in Sec. 1, we believe the success of the monochromatic network may result from three major reasons: 1) eliminating the redundant connections (correlations) between different color channels; 2) reducing the high-dimensional space of RGB input/output by introducing the prior that the rendering process of each color channel (or wavelength) is independent; thus 3) making the network more robust to the inequality and bias in the color distribution of training data. To further verify the superiority of our monochromatic model, we introduce scenes illuminated with colored lights that are not present in our training data (see Fig. 1 and Fig. 13). Our monochromatic model can still produce plausible outputs without further training or fine-tuning, whereas the normal models generally cannot achieve this with no (or only a few) training data in this domain.
@@ 6 Conclusion, Limitation, and Future Works
Acknowledgments
This work was supported by Southern Marine Science and Engineering Guangdong Laboratory (Zhuhai) (No. SML2021SP101) and Shenzhen Science and Technology Program (KJZD20230923114114028).